1.0 Three equal charges of +6 nC are located at the corners of an equilateral triangle (3) whose sides are 12 cm long, as shown below. Find the potential at the center of the triangle.



1
Expert's answer
2022-03-28T14:25:55-0400

Potential

"V_1=V_2=V_3=\frac{kq}{r}"

"V_1=V_2=V_3=\frac{9\times10^9\times6\times10^{-9}\times\sqrt3}{0.12}"

"V_1=V_2=V_3=779.42V"


"V=V_1+V_2+V_3=779.42+779.42+779.42"

"V=2338.26V"